OpenStudy (anonymous):

point/slope form and slope/intercept form for: x-intercept = 6 and y-intercept = -8

OpenStudy (texaschic101):

ok.....your points are (6,0) and (0,-8) find the slope.... m = (y2 - y1) / (x2 - x1) m = (-8 - 0)/(0 - 6) m = -8/-6 = 4/3 <-- this is the slope point slope form is y - y1 = m(x - x1) using 4/3 as slope and point (6,0) y - 0 = 4/3(x - 6) <--point slope form slope intercept form : y = mx + b using 4/3 slope and point (0,-8) -8 = 4/3(0) + b -8 = 0 + b -8 = b y = 4/3x - 8 <--slope intercept form

OpenStudy (anonymous):

ty!!!!!!!!!!!!!!!!!

OpenStudy (texaschic101):

no problem :)